### v1

# MinD-3D: Reconstruct High-quality 3D objects in Human Brain

Abstract
In this paper, we introduce Recon3DMind, a groundbreaking task focused on reconstructing 3D visuals from Functional Magnetic Resonance Imaging (fMRI) signals. This represents a major step forward in cognitive neuroscience and computer vision. To support this task, we present the fMRI-Shape dataset, utilizing 360-degree view videos of 3D objects for comprehensive fMRI signal capture. Containing 55 categories of common objects from daily life, this dataset will bolster future research endeavors. We also propose MinD-3D, a novel and effective three-stage framework that decodes and reconstructs the brain’s 3D visual information from fMRI signals. This method starts by extracting and aggregating features from fMRI frames using a neuro-fusion encoder, then employs a feature bridge diffusion model to generate corresponding visual features, and ultimately recovers the 3D object through a generative transformer decoder. Our experiments demonstrate that this method effectively extracts features that are valid and highly correlated with visual regions of interest (ROIs) in fMRI signals. Notably, it not only reconstructs 3D objects with high semantic relevance and spatial similarity but also significantly deepens our understanding of the human brain’s 3D visual processing capabilities. Project page at: https://jianxgao.github.io/MinD-3D.
Figure 1: We introduce a novel task termed as Recon3DMind. Left: we showcase our data acquisition and modeling principles. Right: our proposed MinD-3D decodes fMRI signals to recover semantically matched, structurally coherent 3D objects with remarkable fidelity.
1 Introduction
Functional Magnetic Resonance Imaging (fMRI), a kind of signal that can be obtained in a non-invasive way, could capture blood changes in the human brain induced by neuronal activity. Due to its relatively easy accessibility, fMRI has been commonly used to reflect visual activities. Some recent studies [4, 5, 31, 27, 26] have successfully reconstructed high-quality images from fMRI signals either by utilizing powerful generative models e.g., diffusion models [29] or using generative transformers [3]. These approaches focus on extracting semantic features from fMRI signals, often requiring only semantic features to generate relevant high-quality images.
However, the human visual system encompasses more than the mere processing of 2D images; it possesses the remarkable ability to translate 2D projections into rich, 3D representations. This complex mechanism bestows upon us a rich perception of our world, complete with discernible aspects of size, distance, and depth. Diverging from earlier studies, our research focuses on modeling this 3D visual capability of the brain. We introduce a novel task, termed Reconstructing 3D Objects from Mind in fMRI (Recon3DMind), that leverages advanced computer vision techniques to decode and reconstruct the 3D visual information perceived by the brain from fMRI signals. This task goes beyond extracting semantic features and includes spatial and structural dimensions, which are crucial for a thorough understanding of 3D visuals.
Some literature [19, 13, 15, 20] has proved that the brain’s visual processing for 3D perception is far more complex than for 2D perception. Consequently, relying solely on semantic features is inadequate for accurately modeling the brain’s capability for 3D spatial imaging. To describe 3D objects effectively, it is essential to consider their shape and structural attributes in addition to their semantic properties. For instance, two cars of varying lengths might look identical from the front but show significant differences when viewed from the side. This observation underscores the importance of capturing the full range of spatial and structural features for a true representation of 3D objects. Therefore, our task focuses on enhancing modeling of human 3D comprehension. It calls for a sophisticated fMRI feature extractor that can discern not just semantic elements but also the spatial structure and other 3D-specific features from fMRI signals, thus facilitating a more complete and accurate recovery of 3D visual information.
Addressing the unique challenges of our Recon3DMind task, we confront a significant hurdle: the absence of a dataset pairing fMRI data with 3D visuals. To fill this void, we introduce fMRI-Shape, a pioneering dataset specifically designed for this purpose. Drawing inspiration from ZeroNVS [30], which posits that 3D objects can be effectively represented through 360-degree view videos, we adopt this methodology for fMRI data collection. In our approach, we present subjects in ShapeNet [2] with 360-degree view videos of stationary 3D objects, where the camera circumnavigates the object, completing a full rotation. This setup enables the capture of fMRI signals as subjects observe the objects from every conceivable angle. Such comprehensive visual exposure ensures the collection of detailed and precise fMRI signals, effectively capturing the entire spectrum of the object’s spatial attributes. Through intricate preprocessing steps, these recordings are transformed into multi-frame fMRI signals. The nuances and intricacies of the fMRI-Shape dataset will be elaborated further in Sec. 3.
With our specially curated fMRI-Shape dataset, we present an innovative and effective three-stage framework to extract both spatial structure and semantic features from multi-frame fMRI signals, named Mind-3D.
In the first stage, we utilize a pre-trained encoder [26] on the NSD [1] dataset to extract spatial features from the fMRI data. These features are then aggregated across multiple frames using a feature aggregation module. To ensure the biological relevance and effectiveness of these features, we align them with the visual-spatial features of the corresponding objects using contrastive learning loss. This alignment guarantees that the features we extract are not only accurate but also meaningfully connected to the brain’s visual processing.
The second stage focuses on decoding the brain’s visual activity features from the extracted fMRI data. We achieve this through a diffusion model trained within the feature space, conditioned on the fMRI features. This conditioning allows us to progressively generate accurate visual representations, effectively translating complex brain activity into comprehensible visual data.
The final stage aims to reconstruct the 3D models as they are perceived by the human brain. To accomplish this, we train a latent-adapted Argus [39] model. This model uses the visual features generated in the second stage along with the initially extracted fMRI features to create 3D objects. This process mirrors the imaging function of the brain’s visual cortex, offering a simulated insight into how the brain perceives and processes 3D objects.
To evaluate the efficacy of our model, we use both semantic and structural metrics. These metrics provide a comprehensive evaluation of the model’s ability to create structurally accurate 3D representations. Additionally, we validate the relevance of our model’s extracted features by correlating them with specific brain regions, ensuring that our model not only generates accurate representations but also does so in a way that is consistent with how the brain processes visual information.
Our contributions can be summarized as follows:
  • We for the first time define the novel task, Recon3DMind, which is meaningful for both computer vision and cognitive neuroscience communities.
  • We introduce an innovative, versatile brain decoding framework, composed of three distinct modules: a neuro-fusion encoder, a feature bridging diffusion model, and a latent adapted decoder.
  • We present fMRI-Shape, the first dataset of fMRI-3D shape pairs, to foster further research in this domain. This dataset will be accessible on our project page.
2 Related work
2.1 FMRI Decoding Methods
Current fMRI decoding methods primarily aim to reconstruct the real world in a 2D format, such as the images or videos that humans perceive. This task is quite complex, as it involves accurately extracting relevant features from fMRI signals and reconstructing a precise 2D representation. Deep learning methods, known for their impressive capabilities, are well-suited for addressing this challenge. Initial successes in this area have been demonstrated by earlier methods [18, 37, 33]. Subsequent studies [32, 7] have shown that generative models are more effective for these tasks. As a result, various diffusion models [4, 5, 31] have been employed as decoders to reconstruct visual scenes, achieving remarkable results. However, these studies have been limited to 2D visual representations. In this paper, we aim to extend the scope of fMRI decoding to 3D representations. Our goal is to directly reconstruct 3D objects from fMRI signals. To accomplish this, we propose a new framework that uses a transformer-based feature encoder for extraction and aggregation. This framework translates neural space data into visual space and employs a powerful 3D decoder to reconstruct the 3D object, utilizing features from the visual space.
2.2 Diffusion Model
Diffusion models [17, 34] are exceptional generative tools for both pixel and feature generation. As a variant, the latent diffusion model [29], equipped with an autoencoder, compresses images into lower-dimensional latent features, thereby generating a compressed version of the data rather than directly generating the data. This approach, which operates in the latent data space, significantly reduces computational requirements and enables the generation of higher-quality images with enhanced details in the latent space. In this paper, we aim to leverage the potent feature generation capabilities of diffusion models to generate visual features based on fMRI features. To achieve this, we have adapted the latent diffusion model by discarding its autoencoder structure and focusing solely on its latent component. The conditional information driving the model is derived from the fMRI features.
2.3 3D Generation
Advancements in deep generative models [36, 41] and 3D-aware image synthesis techniques [6, 38] have significantly influenced the field of 3D object generation. These models are instrumental in creating complex 3D shapes and structures, especially in architecture and virtual environment design, emphasizing discrete voxel generation and 3D modeling from 2D images. In our study, we aim to generate 3D objects from fMRI data, leveraging Argus [40], a robust 3D generative model, as our decoder and adapt it with several transformer layers. This approach facilitates an integrated method that combines fMRI and visual features produced by the previous module. The visual features serve as conditional embeddings for Argus, while the fMRI features are fused in the adapter layers. This synergistic integration aims to enhance the model’s capability to accurately reconstruct 3D objects from complex brain activities.
Figure 2: Overview of the fMRI-Shape Acquisition Process. Initially, we render each object into an 8-second long video, showcasing a 360-degree view. Following this, we capture fMRI signals in video format. The fMRI data are then processed using fMRIPrep, converting the data from 32k fs LR surface space into 2D images with dimensions of 1023 × 2514.
3 Experimental Designs and Curated Dataset
In this section, we will introduce the detailed procedure for the collection of the proposed dataset, fMRI-Shape. This dataset will be released to the community to push the research on the topic of Recon3DMind.
3.1 Subjects and Experiments
Three participants (male, aged 24-27), who were unaware of our work, were instructed to view videos depicting the gradual 360-degree rotation of 3D objects at a 60-degree pitch angle, as shown in Fig. 2. All participants had normal or corrected-to-normal visual acuity. Informed written consent was obtained from each participant, and the experimental protocol was approved by the ethical review board. To ensure the diversity of the dataset for our work, the 3D objects used in this study were derived from the ShapeNetCore [2], which includes 55 classes of objects represented as 3D point clouds, 2D images, and CAD models. We randomly selected four objects from each class in ShapeNetCore, allocating three for training and one for testing. We utilized the rendering technique from Zero123 [21] to generate an 8-second long video for each object in Blender.
For each participant, the experiment was divided into six runs. During each run, participants consecutively viewed 55 videos in a randomized order, one from each category, resulting in an approximate run duration of 8 minutes. The runs alternated between training and test data, with a 1-minute rest period after each run. Notably, each participant underwent three viewings of the testing run and a single viewing of the three training runs.
3.2 Data Acquisition and Preprocessing
MRI data were collected at Zhangjiang International Brain Imaging Centre. T1, T2 and fMRI data were acquired in a 3T scanner and a 32-channel RF head coil. T1-weighted data were scanned using MPRAGE sequence (0.8-mm isotropic resolution, TR=2500ms, TE=2.22ms, flip angle $8^{\circ}$). T2-weighted data were acquired using SPACE sequence 1(0.8-mm isotropic resolution, TR=3200ms, TE=563ms, flip angle $120^{\circ}$). Functional data were scanned using gradient-echo EPI at 2-mm isotropic resolution with whole-brain coverage (TR=800ms, TE=37ms, flip angle $52^{\circ}$, multi-band acceleration factor 8). The sampling frequency of the 3T scanner is 8Hz, so each video corresponds to a total of 10 frames of task-state fMRI signals.
Stimuli were presented using an LCD screen ($8^{\circ}\times 8^{\circ}$) positioned at the head of the scanner bed. Participants viewed the monitor via a mirror mounted on the RF coil and fixated a red central dot (0.4° × 0.4°).
Preprocessing was performed using fMRIPrep ([10, 9]). Following [27], the preprocessed functional data in 32k_fs_LR surface space were converted into 2D images and utilized for further analysis. Specifically, for each 32k_fs_LR data frame in each run, we applied z-scoring to the values across every vertex within the run. These z-scored values were then mapped onto 1023 × 2514 2D images using pycortex [11]. The ROIs for analysis were selected from the HCP-MMP atlas [12] in the 32k_fs_LR space, including “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3”.
4 Problem Setup
Figure 3: Overview of the MinD-3D Framework. Our approach combines a Neuro-Fusion Encoder for extracting features from fMRI frames, a Feature Bridge Diffusion Model for generating visual features from these fMRI signals, and a Latent Adapted Decoder based on the Argus 3D shape generator for reconstructing 3D objects. This integrated system effectively aligns and translates brain signals into accurate 3D visual representations. Note that the CLIP encoder is only for training the model, while it is not used for the inference stage.
Recon3DMind addresses a task of paramount importance in cognitive neuroscience. Developing computational models that accurately interpret and reconstruct the human brain’s 3D comprehension can significantly advance the field of computer vision. This endeavor not only bridges a gap between cognitive neuroscience and computer vision but also promises to propel the latter field forward in unprecedented ways. In this paper, we delve into the specifics of fMRI-based 3D reconstruction, providing detailed definitions and formulas underpinning our methodology.
Consider the acquisition of a multi-frame fMRI signal denoted as $\{F\}$, where $|F|=10$. These signals correspond to both a 3D object point cloud, $\{P\}$, and a video, $\{V\}$, with $|V|=192$ frames, which the subject views. The task necessitates an efficient encoder, $E$, capable of extracting both spatial structural and semantic features from the fMRI signal. It’s crucial to note that while a single frame of the fMRI signal suffices for semantic information extraction in 2D image reconstruction, the reconstruction of 3D structures demands additional spatial structural features. Therefore, we input multiple frames of fMRI signals with the aim of extracting these comprehensive spatial structural features from the spatio-temporal data. This process is mathematically represented as: $f=E(F)$ Following the feature extraction, we employ a powerful decoder to reconstruct the original 3D point cloud $\{P\}$ based on the extracted feature $f$: $P=D(f)$ Consequently, our model can be succinctly defined as $M=\{E,D\}$, thereby operationalizing the transformation: $P=M(F)$.
In executing this model, it’s critical to fully utilize both the fMRI signal $\{F\}$ and the corresponding video $\{V\}$ to train our model $M$. To achieve this, we propose a three-stage innovative and efficient model. Each stage is meticulously designed to capture different aspects of the fMRI data and the associated visual stimuli, ensuring a comprehensive and accurate 3D reconstruction from the complex neural signals. This process not only tests the limits of current computer vision techniques but also offers valuable insights into the human brain’s processing of 3D spatial information.
5 Method
5.1 Overview
As depicted in Fig. 3, our MinD-3D model comprises three main components: 1) Neuro-Fusion Encoder: This part is responsible for extracting both semantic and structural features from the fMRI frames. 2) Feature Bridge Diffusion Model: This module generates visual features based on the features extracted by the Encoder. 3) Latent Adapted Decoder: Finally, this component recovers 3D objects, using both the fMRI features and the generated visual features.
5.2 Neuro-Fusion Encoder
Drawing from the potent Masked Image Model (MIM) [16, 26], we adopt an auto-encoder structure to extract meaningful fMRI features, trained on the UKB dataset [23]. Our encoder, termed the Neuro-Fusion Encoder (NFE), utilizes the encoder of LEA [26] as $E_{f}$ with a Feature Aggregation module $\mathcal{FA}$ to transform fMRI frames into an embedding compatible with the CLIP [28] ViT-B/32 model’s space.
Initially, we process each fMRI frame in parallel to obtain the spatial fMRI embeddings:
$\mathbf{F}_{emb}=E_{f}(\mathbf{F})$
Then, we use the Feature Aggregation module to aggregate the fMRI embeddings into fMRI latent feature $\mathbf{c}_{f}$:
$\mathbf{c}_{f}=\mathcal{FA}(\mathbf{F}_{emb})$
To align the fMRI latent feature with the vision space, we employ the ViT-B/32 CLIP vision encoder $E_{v}$ and here to encode video frames into vision embeddings. During training, we randomly select n frames as $\{\mathbf{V}\}$ and compute the average vision embeddings $V_{emb}$ as the visual latent feature $\mathbf{c}_{v}$:
$\mathbf{c}_{v}=\frac{\sum^{n}_{i=1}E_{v}(\mathbf{V_{i}})}{n}$
where n is the number of selected video frames. Then we utilize the CLIP loss to ensure the alignment between the fMRI features $\mathbf{c}_{f}$ and the visual features $\mathbf{c}_{v}$:
$\mathcal{L}_{c}=\mathcal{L}_{clip}(\mathbf{c}_{f},\mathbf{c}_{v})$
During training, we optimize the whole Neuro-Fusion Encoder with $E_{LEA}$ loading the pre-trained weights and the CLIP vision encoder frozen. It’s worth noting that the CLIP vision encoder $E_{v}$ and images $\{\mathbf{V}\}$ are just used for training, we will drop it during inference. By adopting this approach, we will get the fMRI features $\mathbf{c}_{f}$ and embedding $\mathbf{F}_{emb}$ aligned with the visual space. The fMRI features $\mathbf{c}_{f}$ will be used as the conditional information for the feature bridge diffusion model and fMRI embedding $\mathbf{F}_{emb}$ will be used in the latent adapted decoder.
Figure 4: The qualitative results generated by LEA-3D, fMRI-PTE-3D, and our method are presented. GT indicates the ground-truth 3D objects. All the objects have been rendered into a 2D format.
Methods Semantic-Level Structure-Level
[…]
LEA-3D[26] 0.787 0.371 0.353 0.527 0.562 0.575 0.624 0.372
fMRI-PTE-3D[27] 0.815 0.392 0.375 0.433 0.694 0.327 0.519 0.135
w/o Both 0.784 0.364 0.340 0.536 0.618 0.498 0.649 0.246
w/o Diffusion 0.802 0.382 0.372 0.509 0.641 0.367 0.546 0.175
w/o Contrastive 0.822 0.409 0.398 0.453 0.674 0.236 0.504 0.095
MinD-3D (full) 0.835 0.425 0.411 0.419 0.704 0.214 0.480 0.065
Table 1: Performance Comparison on the fMRI-Shape. We report average metrics for all subjects. LEA-3D and fMRI-PTE-3D represent variants of LEA and fMRI-PTE, respectively. The following 3 baselines are degenerate versions of MinD-3D.
5.3 Feature Bridge Diffusion Model
This phrase details the Neural-Visual Synthesis process using a diffusion model, named here as the Feature Bridging Diffusion Model (FBDM), following the design outlined in [29]. This module’s primary function is to facilitate the generation of the target visual latent feature $\hat{\mathbf{c}_{v}}$ from the corresponding fMRI latent feature $\mathbf{c}_{f}$.
In the forward diffusion process, we treat the visual latent feature $\mathbf{c}_{v}$ as $\mathbf{x}_{0}$ and introduce Gaussian noise into it across 100 timesteps. This approach modifies the standard 1000-timestep model to align with the settings in Mind-eye [31]:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
where, $\mathbf{\epsilon}$ is the Gaussian noise. During the reverse diffusion process, the module will learn to predict and then reverse the noise introduced during the forward process. The goal is to accurately estimate the noise $\hat{\mathbf{\epsilon}}_{t}$ at each timestep, conditioned on the fMRI latent feature $\mathbf{c}_{f}$:
$\hat{\mathbf{\epsilon}}_{t}=FBDM(\mathbf{x}_{t},t,\mathbf{c}_{f})$
The fMRI latent feature $\mathbf{c}_{f}$ is mapped to the intermediate layers of the FBDM through a cross-attention operation:
CrossAttention$(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V$, with
$Q=W_{Q}^{(i)}\varphi_{i}\left(x_{t}\right),K=W_{K}^{(i)}\tau_{\theta}(\mathbf{c}_{f}),V=W_{V}^{(i)}\tau_{\theta}(\mathbf{c}_{f}).$
Here, $\tau_{\theta}$ is the fMRI latent feature with a suitable dimension projector, $\varphi_{i}\left(x_{t}\right)$ denotes intermediate values of the Decoder and $W_{Q}^{(i)}$, $W_{K}^{(i)},W_{V}^{(i)}$ are projector matrices with learnable parameters.
So the training objective of FBDM is to minimize the difference between the actual and predicted noise:
$\mathcal{L}_{FBDM}=\mathbb{E}_{\mathbf{x},\mathbf{\epsilon},t}\left[\|\mathbf{\epsilon}-\hat{\mathbf{\epsilon}}_{t}\|^{2}\right]$
Post-training, the visual latent feature $\mathbf{c}_{v}$ will be dropped and FBDM initiates a generative phase, starting with pure noise and progressively applying the reverse diffusion process to produce the predicted visual latent feature $\hat{\mathbf{c}_{v}}$. This phase crucially maps the aligned fMRI latent features $\mathbf{c}_{f}$ to the visual space and generates the corresponding visual latent features $\hat{\mathbf{c}_{v}}$, which is used as the basic conditional information for the latent adapted decoder.
5.4 Latent Adapted Decoder
Latent Adapted Decoder is based on Argus [39], a cutting-edge 3D shape generator. It will generate the VQ latent representation of the target 3D object based on the conditional information we provide via a generative auto-regressive transformer and then get the 3D objects through the VQ decoder. Here, we design an adapter layer in the transformer block to modify the module using the fMRI embeddings $\mathbf{F}_{emb}$.
Concretely, to directly incorporate the fMRI feature into the 3D generation process, we adapt the Multihead Self-Attention (MSA) in the standard transformer block into a double attention layer. This layer consists of the original MSA and an adapter layer. The adapter layer calculates multi-head cross-attention using the intermediate values of the module. We then obtain the adapted result using a learnable parameter $\lambda$ as the coefficient:
$x=\lambda\cdot\textsc{CA}(Q,K_{f},V_{f})+(1-\lambda)\cdot\textsc{SA}(Q,K,V)$
where CA and SA indicate cross-attention and self-attention separately; $Q,K,V$ denotes the representation of 3d latent features with a Linear projection; $K_{f},V_{f}$ are the representation of fMRI embeddings with a linear projection. The adapter layer is added every T transformer block.
Our decoder computes the VQ latent vector for the target object:
$p(\mathbf{z})=\prod_{i=1}^{m}p\left(\mathbf{z}_{i}\mid\hat{\mathbf{c}_{v}},\mathbf{F}_{emb},\mathbf{z}_{<i}\right)$
Finally, the 3D VQ-decoder decodes this latent vector $\mathbf{z}$ to reconstruct the final point clouds $\mathbf{P}$:
$\{\mathbf{P}\}=D(\mathbf{z})$
The decoder’s optimization objective is defined by the negative log-likelihood loss:
$\mathcal{L}_{nll}=\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[-\log p(\mathbf{z})\right]$
During training, we train the module separately using pre-training weights, after training the neuro-fusion encoder and the feature bridge diffusion model.
6 Expriment
6.1 Metrics
In order to evaluate the performance of the new task and our model comprehensively, we adopt the following metrics in two aspects, semantic-level and structure-level:
Semantic Level. To evaluate the semantic quality of our model, we adopt standard metrics used in previous 2D fMRI literature [24, 4, 5, 25, 22], namely pixel-level correlation (PixCorr) and N-way top-K accuracy. These two metrics are calculated by comparing the reconstructed images with the ground truth (GT) images. In this process, we render both the reconstructed objects and the GT objects into images separately, calculate the metrics for each frame, and then derive the average value as the final metric numbers.
Structure Level. Besides evaluating the semantic quality of our model, it is also essential to assess its performance in terms of spatial structure, which we do in both 3D and 2D. In 3D, we utilize common metrics used in 3D reconstruction: Chamfer Distance (CD), Hausdorff Distance (HD), and Earth Mover’s Distance (EMD). For the 2D evaluation, we render the 3D objects onto a 2D plane and compute the frame-averaged pixel-level correlation, Learned Perceptual Image Patch Similarity (LPIPS), and Structural Similarity (SSIM).
6.2 Implementation Details.
As described in Section 3, each data pair in the fMRI-Shape dataset consists of 10 fMRI frames and the corresponding 192 images. To maximize the utilization of this dataset, we randomly select 6 fMRI frames for training, which also serves as a form of data augmentation. These 6 frames, initially 1023 × 2514 2D fMRI images, are resized to 256 × 256. For the image data, we set the number of selected video frames (n) to 4, meaning we randomly choose 4 frames for subsequent training, and each of these images is resized to 224 × 224. During inference, we select only the middle 6 fMRI frames and do not input any additional images. In all experiments, we train the model using data from all 3 subjects and then calculate the metrics for each subject.
Regarding the model structure, our neuro-fusion encoder comprises 24 transformer layers with 1024 feature dimensions and 16 self-attention heads. Utilizing the CLIP ViT-B/32 model, our Feature Aggregation module maps features to a 512-dimensional space. Consequently, the feature bridge diffusion model is configured with a feature dimension of 512 for cross-attention and 8 self-attention heads, with the number of diffusion channels set to 224. The latent adapted decoder is designed with 32 transformer layers, 3072 feature dimensions, and 16 self-attention heads. We set T=8, which means that we add an adapter layer every 8 transformer blocks.
In terms of model training, the first stage involves optimizing both the neuro-fusion encoder and the feature bridge diffusion model. We initialize the neuro-fusion encoder with pre-trained weights, and the feature bridge diffusion model is trained from scratch. In the second stage, we fix the parameters of the first two stages and focus on optimizing the latent adapted decoder. We use the AdamW optimizer with a learning rate of 5e-5 for all experiments and train the model with a batch size of 128 for all subjects.
6.3 Experimental Results
As the first work to model 3D imaging in the human brain, we compare our comprehensive model, MinD-3D, with several baselines, including adaptations of LEA and fMRI-PTE using the same 3D decoder as in MinD-3D. Tab. 1 presents the averaged metrics at both the Structural and Semantic Levels for all subjects. MinD-3D demonstrates superior performance across all metrics: achieving approximately 83.5% accuracy in 2-way-top-1, 42.5% accuracy in 10-way-top-1, a pixel-level correlation of 0.411, and 0.419 in LPIPS. These results indicate that our model not only generates objects with high semantic accuracy but also excels in structural similarity, outperforming baseline methods.
Qualitative results from LEA-3D, fMRI-PTE-3D, and our method are showcased in Fig. 4. MinD-3D consistently produces 3D objects that are structurally very similar to their real counterparts, while maintaining semantic integrity in most instances. A notable example is the follower vase case, where despite the complexity of its semantics and structure, our model effectively captures and reproduces key features. This underscores the robustness of our model in handling this novel task and its capability to produce faithful realizations.
6.4 Analytical and Ablation Study
We further conduct some analytical and ablation studies to evaluate the effectiveness of our model.
Figure 5: Quality of the extracted features. We map the features to brain regions via ridge regression modeling, with voxel-wise Pearson’s correlation coefficients computed between them. The visual function ROI is highlighted in red, indicating correlations in the range of 0 to 0.8. In contrast, other regions are depicted in blue, representing correlations within a range of 0 to 0.5.
Analysis of feature extracted. To understand our feature extractor from a biological perspective, following [14, 37, 35, 24], we employ a linear encoding model here to project the extracted fMRI feature $c_{f}$ to brain activity. For this purpose, we select subject 1 as our representative. Specifically, we employ ridge regression to learn the mapping relationship between $c_{f}$ and fMRI signals in the training set. We then map $c_{f}$ extracted from the test set to the fMRI space and calculate the Pearson correlation coefficient. As illustrated in Fig. 5, we use red to indicate the positive correlation between visual ROIs. For other regional positive correlations, we use blue. A high positive correlation is evident between $c_{f}$ and visual ROIs, indicating that we have successfully extracted valid features. Additionally, we conducted a similar analysis with the visual latent feature $c_{v}$ and visualized its positive correlations. Notably, the correlation of $c_{f}$ is superior to that of $c_{v}$. Furthermore, we calculated the correlation between $c_{f}$ and $c_{v}$, which is very high. This high correlation suggests that our comparative learning is effective and that the fMRI feature space and image space are well-aligned.
Ablation study of fMRI frames. During training, we randomly select 6 fMRI frames, denoted as $F_{n}$, as the input to our model. To explore the impact of different $F_{n}$ values on our model’s performance, we conducted an assessment across various $F_{n}$ values. We report and compare average results for metrics such as 2-way-top-1, 10-way-top-1, SSIM, and CD, focusing on the performance across three subjects. As shown in Tab. 2, $F_{n}=6$ emerges as the optimal choice. This operation not only acts as data augmentation but also, due to the inherent noise in fMRI signals, choosing the middle portion for inference helps in reducing the introduction of noise.
7 Conclusion
$F_{n}$ Semantic-level Structure-Leve
[…]
1 0.674 0.291 0.472 0.853
3 0.762 0.365 0.604 0.526
6 0.835 0.425 0.704 0.214
8 0.829 0.401 0.702 0.225
10 0.818 0.379 0.684 0.273
Table 2: Ablation results for the fMRI frames. It indicates that $F_{n}=6$ is the best choice for MinD-3D.
In this paper, for the first time, we introduce the task of 3D reconstruction based on fMRI signals, accompanied by a dedicated dataset. Technically, we propose a novel three-stage framework tailored for this groundbreaking task. In the initial stage, our model adeptly extracts features from fMRI frames. Subsequently, in the second stage, it employs a diffusion module to translate these fMRI features into the image domain. The final stage involves the transformation of these image features into 3D models using an advanced 3D generation model. Our experimental results and analyses demonstrate the efficacy of our model in accurately extracting fMRI features and converting them into their corresponding 3D objects. This work not only establishes a new domain in neuroimaging and 3D reconstruction but also opens avenues for further research in understanding and visualizing neural representations in 3D vision.
Appendix A More Implementation Details
To more effectively demonstrate the details of our model and its implementation, this section introduces the necessary preliminary knowledge and provides a detailed description of our neuro-fusion encoder and latent adapted decoder.
A.1 Neuro-Fusion Encoder
Our encoder consists of the encoder from the pre-trained LEA [26] and a feature aggregation module. LEA, a transformer-based auto-encoder, encodes fMRI signals and reconstructs them as depicted in Fig. A.1. Unlike the original Masked Autoencoder (MAE) [16], which only masks some patch-based latent codes, LEA masks all such codes. This enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token during training, thereby facilitating the development of a dense fMRI representation within the [CLS] token.
LEA is trained on the UKB dataset [23]. As shown in Fig. A.1, when LEA is directly applied to our fMRI-Shape dataset, it yields rough reconstructions. However, these reconstructions are not highly precise, primarily due to variations in scanners, subjects, and tasks.
Figure A.1: Architecture of LEA. During training, LEA masks all the patch-based latent codes and enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token.
Given that our input fMRI signals are in video form, to efficiently extract both semantic and structural features from this complex data, we opt to fine-tune the encoder with a small learning rate during the initial stage of training with hyperparameters provided in Tab. A.1.
parameter value parameter value
patch size 16 encoder depth 24
embedding dim 1024 encoder heads 16
mask ratio 1.0 fMRI frames 6
mlp ratio 0.99 image frames 4
Table A.1: Hyperparameters used in neuro-fusion encoder.
A.2 Latent Adapted Decoder
Our decoder is adapted from Argus [39]. The original Argus model enhances the quality and diversity of generated 3D shapes by incorporating tri-plane features as latent representations and a discrete codebook for efficient quantization, leveraging the power of transformers for multi-modal conditional generation. It generates latent codes in an auto-regressive manner, similar to VQGAN [8], based on conditional features. Subsequently, it decodes the quantized vectors into point clouds.
In our latent adapted decoder, we add an adapter layer after every 8 transformer blocks to fuse the fMRI feature. During the second stage of training, we focus on optimizing the parameters in the adapter layers and the learnable parameter $\lambda$. The hyperparameters of our latent adapted decoder are detailed in Tab. A.2.
parameter value parameter value
codebook size 8192 encoder depth 32
codebook dim 512 encoder heads 16
embedding dim 3072 adapter layers 4
mlp ratio 0.99 sequence length 1024
Table A.2: Hyperparameters used in latent adapted decoder.
Appendix B FMRI-Shape Details
To better illustrate our proposed dataset, fMRI-Shape, we provide a detailed description in this section.
Figure B.1: Examples of 360-Degree View Videos. These videos effectively showcase the intricate details, lighting, and shadows of the 3D objects. Presented in a sequence from left to right, they depict the order in which the subjects viewed them. Each video boasts a high resolution of 1024×1024, providing a clear and detailed visual experience.
B.1 More Experimental Details
To clarify the experimental details, we showcase our specially rendered 360-degree view videos. In Fig. B.1, six frames are selected from four videos. The camera in each video encircles the object, completing a full 360-degree clockwise rotation, forming an 8-second video with 192 frames. These high-resolution videos (1024×1024 pixels) vividly capture the intricate details, lighting effects, and shadows of the 3D objects. The arrangement of the videos mirrors the order in which they were viewed by the subjects. Displayed on an LCD screen, these videos enable the subjects to observe and comprehend all facets of the 3D objects. This viewing strategy is crucial, as it aids the subjects’ brains in accurately perceiving the objects, a fundamental aspect for gathering precise and high-quality fMRI signals that correspond to the visual stimuli observed by the subjects.
B.2 More Preprocessing Details
In preparing our data for training, we provide a visualization that clearly shows how we identify the visual Regions of Interest (ROIs) in our data. The preprocessing is applied to all fMRI frames.
Figure B.2: Selection Process of Visual ROIs: On the left, the fMRI is depicted in its original signal form, while on the right, the corresponding area is transformed into a 2D image, demonstrating the conversion process.
Refer to Fig. B.2. The first row displays the raw fMRI signals and their 2D image formation. In the second row, on the left, you can see the appearance of the fMRI signal and the locations of the ROIs. On the right, there is a 2D image illustrating where these ROIs are located within the signal, which has been converted from a format known as 32k_fs_LR.
After identifying these ROIs in the 1023×2514 2D fMRI data, we resize them to 256×256, as shown in Fig. B.3. It’s important to note that our Mind-3D model uses these resized, single-channel 2D fMRI signals as its input.
Figure B.3: Converting ROIs in fMRI Signal
Appendix C More Results
Subject Semantic-Level Structure-Level
[…]
sub-001 0.842 0.424 0.411 0.421 0.704 0.212 0.478 0.069
sub-002 0.836 0.424 0.412 0.420 0.705 0.218 0.483 0.062
sub-003 0.829 0.427 0.410 0.416 0.703 0.212 0.479 0.064
Table C.1: Quantitative results for all subjects.
Qualitative results from different views. In Fig. C.1, we showcase a collection of 2D rendered images of 3D objects by MinD-3D from various viewpoints for subject-1 in fMRI-Shape, closely resembling the ground truth (GT) objects both in semantics and structure.
Quantitative results for all subjects. Our proposed dataset, fmri-shape, includes data from three subjects. To demonstrate the robustness of our model, we present comprehensive quantification results for all three subjects. Tab. C.1 showcases these full metrics, indicating that our model consistently performs well across different subjects. These results underscore the model’s robustness in handling subject variability.
Figure C.1: Qualitative results for subject-1 from various viewpoints.
Appendix D Limitations and Future Works
As the first attempt at the task of Reconstructing 3D Objects from Mind in fMRI (Recon3DMind), our MinD-3D has successfully reconstructed 3D objects that are semantically and structurally similar to the originals. However, the reconstructed 3d objects lack detailed texture, color, and other aspects compared to the original point clouds. To overcome this limitation, our future work will focus on developing methods to more effectively extract texture and color details from 3D objects within fMRI signals. We aim to refine our reconstruction method to create more accurate and detailed replications of the original 3D objects, thereby enhancing the fidelity of the reconstructed models and capturing the complexity and richness of the original designs. Regarding our dataset, we recognize that relying solely on the fMRI-Shape dataset is insufficient for the task’s full scope. A notable challenge is mitigating dataset noise due to the cognitive biases individuals might hold towards certain objects. Therefore, it’s necessary for us to expand our fMRI-Shape dataset by including more 3D objects and collecting data from additional subjects, with the hope of advancing research in our proposed Recon3DMind field.